Birch Company normally produces and sells 45,000 units of RG-6 each month. The selling price is $26 per unit, variable costsBirch Company normally produces and sells 45,000 units of RG-6 each month. The selling price is $26 per unit, variable costsBirch Company normally produces and sells 45,000 units of RG-6 each month. The selling price is $26 per unit, variable costs
Answer #1

Solution 1 & 2:

Sales Price per unit = $26

Variable Cost per Unit = $ 18

Contribution Per unit = $26 - $18 = $8

Net Income if plant continues in strike period of 2 Months:

Net Income = Total Contribution - Fixed manufacturing overhead – Fixed Selling Cost

                      = (20000*8) – (180000*2) – (36000*2) = ($272,000)

Net Income if plant shut down in strike period of 2 months:

Revised fixed manufacturing cost = $180,000 - $41,000 = $139,000 per month

Revised fixed selling cost = $36,000 – ($36,000*9%) = $ 32,760 per month

Start up cost at the end of shut down = $14,000

Net Income = Total Contribution - Fixed manufacturing overhead – Fixed Selling Cost – Startup Cost

                       = 0 – (139000*2) – (32760*2) - 14000 = ($357,520)

                                                               

Hence if plant is closed for 2 months than net income will decrease by $85,520 ($272,000 - $357,520).

As net income is decreasing therefore Birch should not close the plant for 2 months.

Solution 3:

If our loss will remain constant $357,520 at opening of plant and closing of plant then Birch Company will be indifferent. Let x is no of units sold at which Birch company will be indifferent between closing the plant and keeping it open.

It means

Loss at the time of opening the plant = $357,520

8x – (180000*2) – (36000*2) = -357520

8x - 360000 – 72000 = -357520

8x = 74480

X = 9310 units

Hence at 9310 units of sales for 2 months period Birch company be indifferent between closing the plant and keeping it open.